OpenStudy (anonymous):

Find the roots of each polynomial equations. 1. x^3-5x^2+2x+8=0 2. 2x^3+13x+17x-12=0 3. x^3-18x+27=0 4. x^3-2x^2-5x+6=0 5. x^3+x^2-5x+3=0

OpenStudy (anonymous):

well this is a lot of work

OpenStudy (anonymous):

:] just for geniuses like you!

OpenStudy (anonymous):

Dont worry the only reason I am here is for helping..

OpenStudy (anonymous):

lets the first one

OpenStudy (anonymous):

the polynomial is complete, that is to say exponents 3, 2, 1, 0

OpenStudy (anonymous):

then you must get all the dividers of the independent term = 8

OpenStudy (anonymous):

D(8) = (+ or -) 8 , 4, 2, 1

OpenStudy (anonymous):

As you see there are 8 dividers

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

we should replace every one instead of X, if you get ZERO after evaluating we have a first root of the polynomial

OpenStudy (anonymous):

so almost always you begin with 1

OpenStudy (anonymous):

(1)^3 - 5(1)^2 + 2(1) + 8 = 1 - 5 + 2 + 8 = 6 1 is not a root

OpenStudy (anonymous):

then -1 (-1)^3 - 5(-1)^2 + 2(-1) + 8 = -1 -5 -2 + 8 = 0 yes -1 is a root

OpenStudy (anonymous):